Math 46 Homework Assignment (02/22 – 02/26)

Assignment will be collected on 03/05 the Friday after the midterm during the class.

Key notions introduced this week: (1) Method of undetermined coefficients;

(2) Polynomial nonhomogeneous terms;

(3) Exponential nonhomogeneous terms;

(4) Trigonometrical nonhomogeneous terms.

(5) Reduction of Order.

Homework problems:

§5.5. 1, 3, 9, 23, 27. §5.6. 2, 7, 18, 31.